OpenStudy (anonymous):

find the value of cot^2(2tan^(-1)(4/3))

myininaya (myininaya):

\[\cot^2(2 \tan^{-1}(\frac{4}{3}))\] \[\text{ Let } y=\tan^{-1}((\frac{4}{3}))\] => \[\tan(y)=\frac{4}{3}\]  The hyp. is 5. \[\cot^2(2 \cdot y)=[\cot(2y)]^2=[\frac{\cos(2y)}{\sin(2y)}]^2=[\frac{\cos^2(y)-\sin^2(y)}{2 \sin(y) \cos(y)}]^2\] \[=[\frac{[\cos(y)]^2-[\sin(y)]^2}{2 \sin(y) \cos(y)}]^2\] \[=[\frac{[\frac{3}{5}]^2-[\frac{4}{5}]^2}{2 \cdot \frac{4}{5} \cdot \frac{3}{5}}]^2\] \[=[\frac{\frac{9}{25}-\frac{16}{25}}{2 \cdot \frac{12}{25}}]^2\] \[=[\frac{9-16}{2 \cdot 12}]^2\] \[=[\frac{-7}{24}]^2=\frac{49}{24^2}\]

OpenStudy (anonymous):

can you help me with the rest of the problem?

myininaya (myininaya):

What do you mean? Doing 24^2?

myininaya (myininaya):

24^2 means 24*24 You can do the multiplication. I believe in you.

OpenStudy (anonymous):

hahahaha no i know how to do that but the whole problem was sec(tan^-1(-4/3)-csc^-1(-13/12))cot^2(2tan^-1(4/3)

myininaya (myininaya):

It is hard for me to read that I did \[cot^2(2 \cdot \tan^{-1}(\frac{4}{3}))\] Which is the problem I read above? Is that not the right problem?

OpenStudy (anonymous):

yeah you did the last part of the problem the part you have to multiply with the first part

OpenStudy (anonymous):

\[\sec(\tan^-1(-4/3)-\csc^-1(-13/12))\cot^2(2\tan^-1(4/3))\] is that better?

myininaya (myininaya):

\[\sec(\tan^{-1}(\frac{-4}{3})-\csc^{-1}(\frac{-13}{12})) ?\] So you are saying this is the first part?

OpenStudy (anonymous):

yes and what you solved was the second part

myininaya (myininaya):

Which was the only part you had lol But ok...

myininaya (myininaya):

So we have Let y=arctan(-4/3) and let x=arccsc(-13/12) Then we have \[\sec(y-x)=\frac{1}{\cos(y-x)}=\frac{1}{\cos(y)\cos(x)+\sin(y)\sin(x)}\]

myininaya (myininaya):

So draw two right triangles for both of those Let parts...